OpenStudy (anonymous):

simplify sqrt 28a^2 over c^2

OpenStudy (anonymous):

\[\sqrt{28a^2/c^2}\] = 2a/c * \[\sqrt{7}\]

OpenStudy (anonymous):

7√

jhonyy9 (jhonyy9):

(2asqrt7)/c